# ShortcutProbe: Probing Prediction Shortcuts for Learning Robust Models

### v1

@@ 4.2 Prediction Shortcut Detection
$\mathcal{L}_{\text{reg}}=\mathop{\mathbb{E}}\limits_{y\in\mathcal{Y}}\mathop{\mathbb{E}}\limits_{(x,y)\in\mathcal{D}^{y}_{\text{cor}}\cup\mathcal{D}^{y}_{\text{pre}}}\|g_{\psi}(\mathbf{v})-\mathbf{v}\|_{2}^{2}$ (5)
@@ 4.3 Spurious Bias Mitigation
We incorporate the above two loss terms into the overall training objective as follows,
@@ 4.5 Theoretical Analysis
$\gamma_{\varphi}$ $=\mathbb{E}_{\tilde{x},x}\Big{[}\frac{\varphi(\tilde{x})^{T}\mathbf{O}\varphi(x)}{\|\mathbf{O}\varphi(x)\|_{2}^{2}}\Big{]}.=\frac{\mathbb{E}_{\tilde{x}}[\|\mathbf{O}\varphi(\tilde{x})\|_{2}]}{\mathbb{E}_{x}[\|\mathbf{O}\varphi(x)\|_{2}]}$ (13)
@@ 5.1 Datasets
Table 1: Worst-group (WGA) and average accuracy (%) comparison with baseline methods on the Waterbirds, CelebA and Chexpert datasets. Best results are in boldface. All bias mitigation methods use the same half of the validation set.
@@
Table 2: Worst-group (WGA) and average accuracy (%) comparison with baseline methods on the MultiNLI and Civilcomments datasets. Best results are in boldface. All bias mitigation methods use the same half of the validation set.
@@ 5.2 Experimental Setup
We first trained a base model initialized with pretrained weights using empirical risk minimization (ERM) on the training dataset. Then, we retrained the model on half of the validation set using various bias mitigation methods. For our method, we first constructed the probe set using the same half of the validation set and used the probe set for shortcut detection and mitigation. The remaining half of the validation set is used for model selection and hyperparameter tuning. For experiments on the Waterbirds, CelebA, and Chexpert datasets, we used ResNet-50 as the backbone networks, and we used ResNet-18 on the ImageNet-9/A and NICO datasets to ensure a fair comparison with baseline methods. For text datasets, we use a pretrained BERT model Kenton and Toutanova (2019). We trained models with each method for three times using different random seeds and reported the averaged results as well as their standard deviations. Detailed training settings are provided in Appendix [^1] Code is available at https://github.com/gtzheng/ShortcutProbe..
@@
Without group labels, we used the worst-class accuracy Yang et al. (2023) for model selection, which is defined as the worst per-class accuracy on an evaluation set. For performance evaluation on the Waterbirds, CelebA, Chexpert, MultiNLI, and Civilcomments datasets, we adopted the widely accepted metric, worst-group accuracy, which is the lowest accuracy among multiple groups of the test set with each group containing a certain spurious correlation. We also calculated the accuracy gap defined as the standard average accuracy minus the worst-group accuracy to measure the degree of a classifier’s spurious biases. A high worst-group accuracy with a low accuracy gap indicates that the classifier is robust to spurious biases and can fairly predict samples from different groups. We adopted average accuracy for the evaluations on the NICO, ImageNet-9, and ImageNet-A datasets as these datasets are specifically constructed to evaluate the robustness against distributional shifts.
5.3 Analysis of the Probe Set
@@
Our method relies on the probe set for detecting prediction shortcuts and mitigating spurious bias. A good probe set can be used to effectively reveal and mitigate spurious biases in a model, such as those curated with group labels Sagawa et al. (2019). Here, we show that our method, ShortcutProbe, performs effectively without group labels when the probe set is carefully selected from readily available sources, such as the training data and held-out validation data.
@@
To demonstrate, we constructed probe sets using the training data or half of the validation set (with the other half reserved for model selection) as described in Section 5.2. For each data source, we varied the proportion $r$ from 0.1 to 0.5. This adjustment created different versions of the probe set, ranging from those containing only samples with highly confident predictions (small $r$’s) to those including all samples from the selected data source ($r=0.5$).
@@
Fig. 2(a) shows the performance of ShortcutProbe measured by worst-group accuracy (WGA) under different probe sets, while Fig. 2(b) presents the sizes of these probe sets. Compared to ERM models, we observe that reusing the training data for retraining results in minimal improvement on the Waterbirds dataset. This is because all the training data can be correctly predicted by the model, resulting in a probe set that is not informative to learn prediction shortcuts.
@@
Our method demonstrates a significant improvement in robustness to spurious biases by leveraging a relatively small amount of held-out data compared to the size of the reused training data. Additionally, our approach benefits most from samples with high prediction confidence, i.e., when $r$ is small. However, as shown in Fig. 2(a), setting $r$ too small results in an insufficient number of training samples, leading to suboptimal WGA performance. In the following, unless otherwise specified, we use half of the validation set to construct the probe set and treat $r$ as a tunable hyperparameter.
@@ 5.4 Main Results
We focus on a challenging and practical setting where group labels are unavailable in both the training and validation data. This scenario requires detecting and mitigating spurious biases using only the data and models available in a standard ERM training setup. As baselines, we selected state-of-the-art last-layer retraining methods DFR Kirichenko et al. (2023) and AFR Qiu et al. (2023), as well as JTT Liu et al. (2021), which retrains the entire model. For DFR, which typically requires group labels, we used class labels instead. All baseline methods listed in Tables 1 and 2 utilized half of the validation data for training. Similarly, our method employed the same half of the validation data to construct a probe set, derived as a subset of this portion. The remaining half of the validation data was reserved for model selection.
As shown in Tables 1 and 2, our method achieves the highest WGA and the smallest accuracy gap between average accuracy and WGA, indicating its ability to strike a strong balance across different data groups. Methods that achieve high average accuracy, such as DFR on the CelebA dataset, prioritize learning spurious features in the probe set. Although maintaining good predictivity on average, DFR still suffer from the prediction shortcuts, as shown by its low WGA. Our method remains effective on larger backbone networks beyond ResNet-50, such as ResNet-152 and ViT (see Appendix). Unlike other baseline methods, our method uses only a portion of available data for training, highlighting the effectiveness of the probe set in detecting and mitigating spurious biases. Notably, when we applied the same probe set with baseline methods, this led to degraded performance in both WGA and accuracy gap, underscoring the unique advantages of our approach.
@@ 5.5 Ablation Studies
Figure 3: Analyses on how (a) prediction shortcuts as well as their regularization strength $\lambda$, (b) semantic regularization strength $\eta$, and (c) number of base vectors $K$ affect a model’s robustness to spurious biases. We report the worst-group accuracy on the CheXpert dataset.

### v2

@@ 4.2 Prediction Shortcut Detection
$\mathcal{L}_{\text{reg}}=\mathop{\mathbb{E}}\limits_{y\in\mathcal{Y}}\mathop{\mathbb{E}}\limits_{(x,y)\in\mathcal{D}^{y}_{\text{cor}}\cup\mathcal{D}^{y}_{\text{pre}}}\|g_{\psi}(\mathbf{v})-\mathbf{v}\|_{2}^{2}.$ (5)
@@ 4.3 Spurious Bias Mitigation
We combine the two loss terms above to formulate the overall training objective as follows,
@@ 4.5 Theoretical Analysis
$\gamma_{\varphi}$ $=\mathbb{E}_{\tilde{x},x}\Big{[}\frac{\varphi(\tilde{x})^{T}\mathbf{O}\varphi(x)}{\|\mathbf{O}\varphi(x)\|_{2}^{2}}\Big{]}=\frac{\mathbb{E}_{\tilde{x}}[\|\mathbf{O}\varphi(\tilde{x})\|_{2}]}{\mathbb{E}_{x}[\|\mathbf{O}\varphi(x)\|_{2}]}.$ (13)
@@ 5.1 Datasets
Table 1: Comparison of worst-group accuracy (WGA) and average accuracy (%) with baseline methods on the Waterbirds, CelebA, and CheXpert datasets. The best results are highlighted in boldface. All bias mitigation methods use the same half of the validation set.
@@
Table 2: Comparison of worst-group accuracy (WGA) and average accuracy (%) with baseline methods on the MultiNLI and Civilcomments datasets. Best results are highlighted in boldface. All bias mitigation methods use the same half of the validation set.
@@ 5.2 Experimental Setup
We first trained a base model initialized with pretrained weights using empirical risk minimization (ERM) on the training dataset. Then, we retrained the model on half of the validation set using various bias mitigation methods. For our method, we first constructed the probe set using the same half of the validation set and used the probe set for shortcut detection and mitigation. The remaining half of the validation set was used for model selection and hyperparameter tuning. For experiments on the Waterbirds, CelebA, and CheXpert datasets, we used ResNet-50 as the backbone network, and we used ResNet-18 on the ImageNet-9/A and NICO datasets to ensure a fair comparison with baseline methods. For text datasets, we used a pretrained BERT model Kenton and Toutanova (2019). We trained models using each method three times with different random seeds and reported the average results as well as their standard deviations. Detailed training settings are provided in Appendix [^1] Code is available at https://github.com/gtzheng/ShortcutProbe..
@@
Without group labels, we used the worst-class accuracy Yang et al. (2023) for model selection, which is defined as the worst per-class accuracy on an evaluation set. For performance evaluation on the Waterbirds, CelebA, CheXpert, MultiNLI, and Civilcomments datasets, we adopted the widely accepted metric, worst-group accuracy, which is the lowest accuracy across multiple groups of the test set with each group containing a certain spurious correlation. We also calculated the accuracy gap defined as the standard average accuracy minus the worst-group accuracy to measure the degree of a classifier’s spurious biases. A high worst-group accuracy with a low accuracy gap indicates that the classifier is robust to spurious biases and can fairly predict samples from different groups. We adopted average accuracy for the evaluations on the NICO, ImageNet-9, and ImageNet-A datasets as these datasets are specifically constructed to evaluate the robustness against distributional shifts.
5.3 Analysis of Probe Set
@@
Our method relies on a probe set for detecting prediction shortcuts and mitigating spurious bias. A good probe set can be used to effectively reveal and mitigate spurious biases in a model, such as those curated with group labels Sagawa et al. (2019). Here, we show that our method, ShortcutProbe, performs effectively without group labels when a probe set is carefully selected from readily available sources, such as the training data and held-out validation data.
@@
To demonstrate, we constructed a probe set as described in Section 5.2, using the training set or half of the validation set (with the other half reserved for model selection) as the data source. For each data source, we varied the proportion $r$from 20% to 100%. This adjustment created different probe sets, ranging from those containing only samples with highly confident predictions (small $r$) to those including all samples from the selected data source ($r=100$).
@@
Fig. 2(a) shows the performance of ShortcutProbe measured by worst-group accuracy (WGA) under different probe sets, while Fig. 2(b) presents the sizes of these probe sets. Compared to ERM models, we observe that using the training data for retraining results in minimal improvement on the Waterbirds dataset. This is because most of the training data can be correctly predicted by the model, resulting in a probe set that is not informative for learning prediction shortcuts.
@@
In contrast, by leveraging a relatively small amount of the held-out data compared to the size of the training data, our method demonstrates a significant improvement in robustness to spurious biases. Additionally, our approach benefits most from samples with high prediction confidence, i.e., when $r$ is small. However, as shown in Fig. 2(a), setting $r$ too small results in an insufficient number of training samples, leading to suboptimal WGA performance. In the following, unless otherwise specified, we use half of the validation set to construct the probe set and treat $r$ as a tunable hyperparameter.
@@ 5.4 Main Results
We focus on a challenging and practical setting where group labels are unavailable in both training and validation data. This scenario requires detecting and mitigating spurious biases using only the available data and model in a standard ERM training setup. As baselines, we selected state-of-the-art last-layer retraining methods DFR Kirichenko et al. (2023) and AFR Qiu et al. (2023), as well as JTT Liu et al. (2021), which retrains the entire model. For DFR, which typically requires group labels, we used class labels instead. All baseline methods listed in Tables 1 and 2 utilized half of the validation data for training. Similarly, we used the same half of the validation data to construct a probe set, derived as a subset of this portion. The remaining half of the validation data was reserved for model selection.
As shown in Tables 1 and 2, our method achieves the highest WGA and the smallest accuracy gap between average accuracy and WGA, indicating its effectiveness in balancing performance across different data groups. Methods that achieve high average accuracy, such as DFR on the CelebA dataset, prioritize learning spurious features in the probe set. Although maintaining good predictivity on average, DFR still suffers from prediction shortcuts, as shown by its low WGA. Our method remains effective on larger backbone networks beyond ResNet-50, such as ResNet-152 and ViT (see Appendix). Unlike other baseline methods, our method uses only a portion of available data for training, highlighting the effectiveness of the probe set in detecting and mitigating spurious biases. Notably, when we applied the same probe set with baseline methods, this led to degraded performance in both WGA and accuracy gap, underscoring the unique advantages of our approach.
@@ 5.5 Ablation Studies
Figure 3: Analyses on how (a) prediction shortcuts as well as their regularization strength $\lambda$, (b) semantic regularization strength $\eta$, and (c) number of base vectors $K$ affect a model’s robustness to spurious bias. We report the worst-group accuracy on the CheXpert dataset.